# MAATS: A Multi-Agent Automated Translation System Based on MQM Evaluation

### v1

@@ 4 Discussion
{CJK*}
UTF8gbsn Beyond the primary evaluation, we conducted a qualitative meta-analysis of three case studies to compare different approaches. We revealed that MAATS consistently excels in three key areas: (1) multi-layered error analysis beyond surface issues, (2) detects omissions and subtleties across perspectives, and (3) enriches context to match target locale expectations. Shown in Fig. 6, MAATS flagged the unnatural literal translation “预定票价” and revised it to the idiomatic “预定机票.” It also uncovered cultural implications in “Car Guy Car” (see Table 3), inferred implied actions in “I started a new shawl” (Table 5), and adjusted tone in “literature class hatred” (Table 6) for classroom context. These examples demonstrate MAATS’s superior sensitivity to context and nuance (full comparisons in Appendix F).
@@ 5 Limitations
{CJK*}
UTF8gbsn While MAATS shows strong performance in error detection and correction, our meta-analysis reveals key limitations shared by both MAATS and Single-Agent systems, particularly in capturing emotional tone, pragmatic nuance, and rhetorical structure. These challenges stem from both model behavior and inherent constraints in the MQM framework. For instance, emotionally charged terms like “hatred” were mistranslated with excessive intensity “仇恨” (see Table 6), missing the speaker’s intended tone. Sentences with emotional contrast, such as “Abe keeps his sorrows hidden and laughs easily,” were misinterpreted as lighthearted, failing to capture the ironic or concealed emotion. Moreover, rhetorical structures like repetition (e.g., “ads, ads, ads”) were either rigidly translated or replaced with fluent idioms, losing the original stylistic emphasis. These examples highlight that while MAATS excels in accuracy and fluency, it struggles with affective meaning conveyed through tone, irony, or form. Crucially, MQM lacks dimensions for evaluating emotional fidelity, speaker stance, or discourse-level intent, leaving such nuances under-annotated and uncorrected. Addressing these issues may require expanding MQM or integrating it with complementary methods like discourse modeling and affective computing to better reflect the full communicative intent in translation.
6 Acknowledgements
I am deeply and sincerely grateful to Professor Safinah Ali, whose support has been foundational to this project from start to finish. Her mentorship went far beyond academic guidance—she generously provided research resources and constant encouragement at every step. This work would not have been possible without her trust and backing.
I would also like to thank Professor Tanya Goyal for taking the time to engage with my work and offer constructive, thoughtful feedback. Her insights during my presentation helped clarify the direction of my analysis.
Finally, my heartfelt thanks go to Shirley—my fiancée and my closest collaborator. She stood by my side through every challenge, offering not only critical contributions to the research but also the emotional strength I needed to keep going. Her presence, patience, and unwavering support carried me through to the very end.
@@ F.2 Case Study 2: “Book the Fare” – MAATS Detects Omissions Across Perspectives
{CJK*}
UTF8gbsn It’s worth noting that the terminology annotation itself was not fully accurate, as it emphasized highlighting “fare” as a term but did not correct its misinterpretation as “票价.” This introduced a degree of internal conflict among annotations, yet MAATS still arrived at the correct fix—translating the phrase as “预定机票.” This case shows that MAATS not only detects more errors but also demonstrates robust decision-making in the face of conflicting signals. The system’s redundancy—flagging the issue multiple times from different perspectives—may serve as a self-correcting mechanism, compensating for occasional misclassifications.
@@ F.3 Case Study 3
{CJK*}
UTF8gbsn The third case, Table 5 and Table 6, highlights MAATS’s ability to enrich translations contextually based on cultural and linguistic expectations of the target locale. The phrase “I started a new shawl” was literally translated with “开始” (to start), producing a grammatically acceptable but culturally unnatural expression in Chinese. A more idiomatic translation would be “我开始织” (I started knitting) or “我开始制作” (I started making), which conveys the intended meaning more clearly in the target language.
@@
{CJK*}
UTF8gbsn Item Single Agent Multi-Agent Source Sentence I was looking for a good first car guy car; and I really love the 80s aesthetic. Same as Single Agent Initial Translation 我在寻找一辆好的第一辆车,而且我真的很喜欢80年代的美学。 Same as Single Agent Accuracy [Major] Mistranslation – Failed to convey "car guy car" as a car for enthusiasts. [Critical] Omission – "Car guy" omitted.
@@
Fluency [Minor] Punctuation – Semicolon translated as comma. [Minor] Punctuation – Semicolon should be a comma in Chinese (“车;而且”). Style [Major] Awkward – Literal rendering of "car guy car" is unnatural. [Minor] Unnatural Flow – “好的第一辆车” sounds awkward.
@@
Audience Appropriateness [Major] Culture-specific Reference – "Car guy car" may not be understood by Chinese readers; adaptation needed. Terminology [Critical] Term Not Applied – "Car guy" was not translated per glossary. Locale Conventions [Minor] Number Format – “第一辆车” doesn’t reflect enthusiast nuance. Other Categories No issues marked in: Linguistic Conventions (other than punctuation), Non-Translation, and Other. Same as left. Final Translation 我在寻找一辆适合车迷的第一辆车;而且我真的很喜欢80年代的美学。 我在寻找一辆适合车迷的新手车,而且我真的很喜欢80年代风格。
@@
{CJK*}
UTF8gbsn Item Single Agent Multi-Agent Source Sentence We provide links to where you can book the fare. Same as Single Agent Initial Translation 我们提供可以预订票价的链接。 Same as Single Agent Accuracy [Major] Omission – The phrase “to where” is omitted, which specifies the destination or platform. [Critical] Omission – “to where” is omitted from the translation.
@@
Fluency [Minor] Punctuation – Missing space after the period in the source. [Minor] Punctuation – Missing period at the end of the Chinese sentence. Style [Minor] Unnatural Flow – “可以预订票价的链接” is correct but slightly awkward. More natural: “可以预订机票的链接” or “票价预订链接”. Audience Appropriateness [Minor] Culture-specific Reference – “预订票价” is not natural in Chinese. Suggested alternatives: “预订机票” or “预订航班”. Terminology [Critical] Term Not Applied – “Fare” should be highlighted per glossary but wasn’t. Locale Conventions No error. Other Categories No issues marked in: Linguistic Conventions (other than punctuation), Non-Translation, and Other. Same as left. Final Translation 我们提供链接,您可以通过这些链接预订票价。 我们提供可以预订机票的链接。
@@
{CJK*}
UTF8gbsn Item Single Agent Multi-Agent Source Sentence I started a new shawl. Same as Single Agent Initial Translation 我开始了一条新的披肩。 Same as Single Agent Accuracy [Minor] Addition – The word “一条” was added in the translation. While it is not present in the source, it does not significantly affect meaning. Fluency No error. Style [Minor] Unnatural Flow – “我开始了一条新的披肩” is grammatically correct, but a more natural Chinese expression is “我开始织一条新的披肩”. Audience Appropriateness [Minor] Culture-specific Reference – “披肩” is appropriate and understandable; no adaptation needed. Terminology No error. Locale Conventions No error. Other Categories No issues marked in any other dimensions. Same as left. Final Translation 我开始了一条新的披肩。 我开始织一条新的披肩。
@@
{CJK*}
UTF8gbsn Item Single Agent Multi-Agent Source Sentence The literature class hatred is even worse: I love to read. Same as Single Agent Initial Translation 文学课的仇恨更严重:我喜欢阅读。 Same as Single Agent Accuracy [Major] Mistranslation – The phrase “The literature class hatred” is misinterpreted and suggests something more severe than intended. [Major] Mistranslation – “仇恨” correctly means “hatred,” but fails to reflect the tone/context of dislike within an academic subject. Fluency [Minor] Punctuation – The colon “:” may not be the best fit for Chinese sentence structure. [Minor] Punctuation – The colon should be a full-width colon “:” to match Chinese conventions. Style [Major] Awkward – The structure is grammatically acceptable but feels unnatural in Chinese. [Major] Unnatural Flow – “文学课的仇恨更严重” is awkward; better options include “我对文学课的厌恶更深” or “反感更强”. Audience Appropriateness [Major] Culture-specific Reference – The phrase “class hatred” may not culturally align with Chinese interpretation of classroom sentiments. Terminology No error. Locale Conventions No error. Other Categories No issues marked in any other dimensions. Same as left. Final Translation 文学课上的仇恨更糟糕,我却喜欢阅读。 文学课的厌恶更深:我喜欢阅读。
@@ Appendix H MAATS Evaluator Prompts
You are responsible for evaluating the audience appropriateness of machine-translated content. Your main task is to detect audience appropriateness issues and judge how severely they affect the reader’s experience. Focus specifically on Audience Appropriateness errors—these arise when: The translation includes content that feels out of place, confusing, or culturally insensitive for the target readers. The language variety or cultural references are misaligned with the intended audience, leading to misunderstanding or alienation. Audience Appropriateness errors are categorized as follows: 1. Culture-Specific Reference: A metaphor, idiom, term, or expression is used that does not fit the audience’s cultural context, making the message unclear or inappropriate. 2. Wrong Language Variety: The translation uses a different regional dialect, spelling, or phrasing than what the audience expects (e.g., UK English instead of US English).   Severity Levels Critical: Severely hinders understanding or creates major cultural offense or disconnect. Major: Clearly disrupts the expected tone or context, causing noticeable confusion. Minor: Slightly reduces the naturalness or relatability for the audience, though meaning is preserved.   {CJK*} UTF8min Example Source (English): Please contact our support team if you experience any issues. Target (Japanese): お客様ごとに担当者が異なりますので、まずは担当者にご相談ください MQM annotations: major: audience_appropriate/culture_specific_reference – The phrase “お客様ごとに担当者が異なりますので…” (“each customer has a different representative”) assumes a Japan-specific corporate support structure that may not reflect the source content. It introduces a culturally specific reference not present in the source and could mislead or confuse international audiences.   Output Format Learn from the example, based on the source segment and the machine translation provided, identify the error types in the translation and classify them using the following template: MQM annotations: Audience Appropriateness Errors [Critical]: [audience_appropriateness/error_subcategory] - Brief explanation. [Major]: [audience_appropriateness/error_subcategory] - Brief explanation. [Minor]: [audience_appropriateness/error_subcategory] - Brief explanation. Please follow the template, put [Error Severity Levels]:[audience_appropriateness/error_subcategory] - None If doesn’t exit
@@
You are responsible for evaluating the design and markup integrity of machine-translated content. Your main task is to detect design or formatting issues and judge how severely they impact the clarity, layout, or usability of the translation. Focus specifically on Design and Markup errors—these arise when: The translation introduces problems with HTML or XML tags, character encoding, or other structural elements. Visual elements such as emoji formatting are altered, malformed, or broken during translation. Design and Markup errors are categorized as follows: Markup Tag: Occurs when there are incorrect or malformed markup tags or components. This includes: Improperly escaped HTML entities (e.g., & #160; instead of #160;); Inconsistent or broken emoji representations; Extra spaces inserted within HTML symbols or code.   Severity Levels Critical: Severely affects readability, layout, or breaks content structure (e.g., broken HTML tags that cause rendering failures). Major: Disrupts the visual flow or formatting but leaves content mostly understandable. Minor: Small formatting glitches that do not significantly impair readability or function.   {CJK*}UTF8gbsn Example
@@
You are responsible for evaluating the locale-specific formatting and conventions of machine-translated content. Your task is to identify issues where the translation fails to follow regional standards or client-specific formatting requirements. Focus specifically on Locale Conventions errors—these arise when: The translation includes content that does not match regional formats for things like dates, numbers, addresses, or telephone numbers. Client instructions for localized formatting are ignored, even if the text is grammatically correct. Locale Conventions errors are categorized as follows: 1. Address Format: The structure of an address does not follow the local convention (e.g., writing city before postal code in a region where the reverse is standard). 2. Currency Format: Currency symbols, abbreviations, or placement are incorrect (e.g., using “\$100” instead of “100€” for European locales). 3. Date/Time Format: Date or time expressions do not align with the regional norm (e.g., “03/04/2023” for UK should mean 3 April, not March 4). 4. Measurement Format: Use of measurement units (e.g., inches, meters, grams) that are inappropriate for the locale or formatted incorrectly (e.g., “5ft” instead of “1.52m”). 5. Number Format: Digits, separators, or groupings deviate from locale rules (e.g., “1,000.00” in US vs. “1.000,00” in many EU countries). 6. Telephone Format: Phone numbers do not follow regional presentation (e.g., missing country code or incorrect spacing).   {CJK*}UTF8gbsn Example
@@
You are responsible for evaluating the correct use of glossary terms in machine-translated content. Your task is to identify issues where required terminology is not used correctly or fails to match the glossary specification. Focus specifically on Terminology errors—these arise when: A glossary term (highlighted in blue in the Annotation Tool) is missing, misapplied, misspelled, or grammatically incorrect. The translation includes the correct glossary term but uses it in a contextually inappropriate way. Terminology errors are categorized as follows: 1. Term Not Applied: A required glossary term was not used; a different or incorrect term was inserted instead, violating glossary rules. 2. Wrong Term: The glossary term appears but is used incorrectly in context—this includes typos, capitalization issues, plural/singular mismatches, or improper grammatical inflection.   Severity Levels Critical: Glossary misuse leads to confusion, misinterpretation, or disrupts key meaning. Major: The error affects tone or grammatical flow, though the meaning remains somewhat clear. Minor: Cosmetic or minor issues such as case sensitivity or small typos that don’t impact comprehension.   {CJK*} UTF8gbsn Example Source (English): S/N (Serial Number) Target (Russian): серийный номер (серийный номер)

### v2

@@ 4 Discussion
Beyond the primary evaluation, we conducted a qualitative meta-analysis of three case studies to compare different approaches. We revealed that MAATS consistently excels in three key areas: (1) multi-layered error analysis beyond surface issues, (2) detects omissions and subtleties across perspectives, and (3) enriches context to match target locale expectations. Shown in Fig. 6, MAATS flagged the unnatural literal translation “预定票价” and revised it to the idiomatic “预定机票.” It also uncovered cultural implications in “Car Guy Car” (see Table 3), inferred implied actions in “I started a new shawl” (Table 5), and adjusted tone in “literature class hatred” (Table 6) for classroom context. These examples demonstrate MAATS’s superior sensitivity to context and nuance (full comparisons in Appendix F).
@@ 5 Limitations
While MAATS shows strong performance in error detection and correction, our meta-analysis reveals key limitations shared by both MAATS and Single-Agent systems, particularly in capturing emotional tone, pragmatic nuance, and rhetorical structure. These challenges stem from both model behavior and inherent constraints in the MQM framework. For instance, emotionally charged terms like “hatred” were mistranslated with excessive intensity “仇恨” (see Table 6), missing the speaker’s intended tone. Sentences with emotional contrast, such as “Abe keeps his sorrows hidden and laughs easily,” were misinterpreted as lighthearted, failing to capture the ironic or concealed emotion. Moreover, rhetorical structures like repetition (e.g., “ads, ads, ads”) were either rigidly translated or replaced with fluent idioms, losing the original stylistic emphasis. These examples highlight that while MAATS excels in accuracy and fluency, it struggles with affective meaning conveyed through tone, irony, or form. Crucially, MQM lacks dimensions for evaluating emotional fidelity, speaker stance, or discourse-level intent, leaving such nuances under-annotated and uncorrected. Addressing these issues may require expanding MQM or integrating it with complementary methods like discourse modeling and affective computing to better reflect the full communicative intent in translation.
@@ F.2 Case Study 2: “Book the Fare” – MAATS Detects Omissions Across Perspectives
It’s worth noting that the terminology annotation itself was not fully accurate, as it emphasized highlighting “fare” as a term but did not correct its misinterpretation as “票价.” This introduced a degree of internal conflict among annotations, yet MAATS still arrived at the correct fix—translating the phrase as “预定机票.” This case shows that MAATS not only detects more errors but also demonstrates robust decision-making in the face of conflicting signals. The system’s redundancy—flagging the issue multiple times from different perspectives—may serve as a self-correcting mechanism, compensating for occasional misclassifications.
@@ F.3 Case Study 3
The third case, Table 5 and Table 6, highlights MAATS’s ability to enrich translations contextually based on cultural and linguistic expectations of the target locale. The phrase “I started a new shawl” was literally translated with “开始” (to start), producing a grammatically acceptable but culturally unnatural expression in Chinese. A more idiomatic translation would be “我开始织” (I started knitting) or “我开始制作” (I started making), which conveys the intended meaning more clearly in the target language.
@@
Item Single Agent Multi-Agent
Source Sentence I was looking for a good first car guy car; and I really love the 80s aesthetic. Same as Single Agent
Initial Translation 我在寻找一辆好的第一辆车,而且我真的很喜欢80年代的美学。 Same as Single Agent
Accuracy [Major] Mistranslation – Failed to convey "car guy car" as a car for enthusiasts. [Critical] Omission – "Car guy" omitted.
@@
Fluency [Minor] Punctuation – Semicolon translated as comma. [Minor] Punctuation – Semicolon should be a comma in Chinese (“车;而且”).
Style [Major] Awkward – Literal rendering of "car guy car" is unnatural. [Minor] Unnatural Flow – “好的第一辆车” sounds awkward.
@@
Audience Appropriateness [Major] Culture-specific Reference – "Car guy car" may not be understood by Chinese readers; adaptation needed.
Terminology [Critical] Term Not Applied – "Car guy" was not translated per glossary.
Locale Conventions [Minor] Number Format – “第一辆车” doesn’t reflect enthusiast nuance.
Other Categories No issues marked in: Linguistic Conventions (other than punctuation), Non-Translation, and Other. Same as left.
Final Translation 我在寻找一辆适合车迷的第一辆车;而且我真的很喜欢80年代的美学。 我在寻找一辆适合车迷的新手车,而且我真的很喜欢80年代风格。
@@
Item Single Agent Multi-Agent
Source Sentence We provide links to where you can book the fare. Same as Single Agent
Initial Translation 我们提供可以预订票价的链接。 Same as Single Agent
Accuracy [Major] Omission – The phrase “to where” is omitted, which specifies the destination or platform. [Critical] Omission – “to where” is omitted from the translation.
@@
Fluency [Minor] Punctuation – Missing space after the period in the source. [Minor] Punctuation – Missing period at the end of the Chinese sentence.
Style [Minor] Unnatural Flow – “可以预订票价的链接” is correct but slightly awkward. More natural: “可以预订机票的链接” or “票价预订链接”.
Audience Appropriateness [Minor] Culture-specific Reference – “预订票价” is not natural in Chinese. Suggested alternatives: “预订机票” or “预订航班”.
Terminology [Critical] Term Not Applied – “Fare” should be highlighted per glossary but wasn’t.
Locale Conventions No error.
Other Categories No issues marked in: Linguistic Conventions (other than punctuation), Non-Translation, and Other. Same as left.
Final Translation 我们提供链接,您可以通过这些链接预订票价。 我们提供可以预订机票的链接。
@@
Item Single Agent Multi-Agent
Source Sentence I started a new shawl. Same as Single Agent
Initial Translation 我开始了一条新的披肩。 Same as Single Agent
Accuracy [Minor] Addition – The word “一条” was added in the translation. While it is not present in the source, it does not significantly affect meaning.
Fluency No error.
Style [Minor] Unnatural Flow – “我开始了一条新的披肩” is grammatically correct, but a more natural Chinese expression is “我开始织一条新的披肩”.
Audience Appropriateness [Minor] Culture-specific Reference – “披肩” is appropriate and understandable; no adaptation needed.
Terminology No error.
Locale Conventions No error.
Other Categories No issues marked in any other dimensions. Same as left.
Final Translation 我开始了一条新的披肩。 我开始织一条新的披肩。
@@
Item Single Agent Multi-Agent
Source Sentence The literature class hatred is even worse: I love to read. Same as Single Agent
Initial Translation 文学课的仇恨更严重:我喜欢阅读。 Same as Single Agent
Accuracy [Major] Mistranslation – The phrase “The literature class hatred” is misinterpreted and suggests something more severe than intended. [Major] Mistranslation – “仇恨” correctly means “hatred,” but fails to reflect the tone/context of dislike within an academic subject.
Fluency [Minor] Punctuation – The colon “:” may not be the best fit for Chinese sentence structure. [Minor] Punctuation – The colon should be a full-width colon “:” to match Chinese conventions.
Style [Major] Awkward – The structure is grammatically acceptable but feels unnatural in Chinese. [Major] Unnatural Flow – “文学课的仇恨更严重” is awkward; better options include “我对文学课的厌恶更深” or “反感更强”.
Audience Appropriateness [Major] Culture-specific Reference – The phrase “class hatred” may not culturally align with Chinese interpretation of classroom sentiments.
Terminology No error.
Locale Conventions No error.
Other Categories No issues marked in any other dimensions. Same as left.
Final Translation 文学课上的仇恨更糟糕,我却喜欢阅读。 文学课的厌恶更深:我喜欢阅读。
@@ Appendix H MAATS Evaluator Prompts
You are responsible for evaluating the audience appropriateness of machine-translated content. Your main task is to detect audience appropriateness issues and judge how severely they affect the reader’s experience. Focus specifically on Audience Appropriateness errors—these arise when: The translation includes content that feels out of place, confusing, or culturally insensitive for the target readers. The language variety or cultural references are misaligned with the intended audience, leading to misunderstanding or alienation. Audience Appropriateness errors are categorized as follows: 1. Culture-Specific Reference: A metaphor, idiom, term, or expression is used that does not fit the audience’s cultural context, making the message unclear or inappropriate. 2. Wrong Language Variety: The translation uses a different regional dialect, spelling, or phrasing than what the audience expects (e.g., UK English instead of US English).   Severity Levels Critical: Severely hinders understanding or creates major cultural offense or disconnect. Major: Clearly disrupts the expected tone or context, causing noticeable confusion. Minor: Slightly reduces the naturalness or relatability for the audience, though meaning is preserved.   Example Source (English): Please contact our support team if you experience any issues. Target (Japanese): お客様ごとに担当者が異なりますので、まずは担当者にご相談ください MQM annotations: major: audience_appropriate/culture_specific_reference – The phrase “お客様ごとに担当者が異なりますので…” (“each customer has a different representative”) assumes a Japan-specific corporate support structure that may not reflect the source content. It introduces a culturally specific reference not present in the source and could mislead or confuse international audiences.   Output Format Learn from the example, based on the source segment and the machine translation provided, identify the error types in the translation and classify them using the following template: MQM annotations: Audience Appropriateness Errors [Critical]: [audience_appropriateness/error_subcategory] - Brief explanation. [Major]: [audience_appropriateness/error_subcategory] - Brief explanation. [Minor]: [audience_appropriateness/error_subcategory] - Brief explanation. Please follow the template, put [Error Severity Levels]:[audience_appropriateness/error_subcategory] - None If doesn’t exit
@@
You are responsible for evaluating the design and markup integrity of machine-translated content. Your main task is to detect design or formatting issues and judge how severely they impact the clarity, layout, or usability of the translation. Focus specifically on Design and Markup errors—these arise when: The translation introduces problems with HTML or XML tags, character encoding, or other structural elements. Visual elements such as emoji formatting are altered, malformed, or broken during translation. Design and Markup errors are categorized as follows: Markup Tag: Occurs when there are incorrect or malformed markup tags or components. This includes: Improperly escaped HTML entities (e.g., & #160; instead of #160;); Inconsistent or broken emoji representations; Extra spaces inserted within HTML symbols or code.   Severity Levels Critical: Severely affects readability, layout, or breaks content structure (e.g., broken HTML tags that cause rendering failures). Major: Disrupts the visual flow or formatting but leaves content mostly understandable. Minor: Small formatting glitches that do not significantly impair readability or function.   Example
@@
You are responsible for evaluating the locale-specific formatting and conventions of machine-translated content. Your task is to identify issues where the translation fails to follow regional standards or client-specific formatting requirements. Focus specifically on Locale Conventions errors—these arise when: The translation includes content that does not match regional formats for things like dates, numbers, addresses, or telephone numbers. Client instructions for localized formatting are ignored, even if the text is grammatically correct. Locale Conventions errors are categorized as follows: 1. Address Format: The structure of an address does not follow the local convention (e.g., writing city before postal code in a region where the reverse is standard). 2. Currency Format: Currency symbols, abbreviations, or placement are incorrect (e.g., using “\$100” instead of “100€” for European locales). 3. Date/Time Format: Date or time expressions do not align with the regional norm (e.g., “03/04/2023” for UK should mean 3 April, not March 4). 4. Measurement Format: Use of measurement units (e.g., inches, meters, grams) that are inappropriate for the locale or formatted incorrectly (e.g., “5ft” instead of “1.52m”). 5. Number Format: Digits, separators, or groupings deviate from locale rules (e.g., “1,000.00” in US vs. “1.000,00” in many EU countries). 6. Telephone Format: Phone numbers do not follow regional presentation (e.g., missing country code or incorrect spacing).   Example
@@
You are responsible for evaluating the correct use of glossary terms in machine-translated content. Your task is to identify issues where required terminology is not used correctly or fails to match the glossary specification. Focus specifically on Terminology errors—these arise when: A glossary term (highlighted in blue in the Annotation Tool) is missing, misapplied, misspelled, or grammatically incorrect. The translation includes the correct glossary term but uses it in a contextually inappropriate way. Terminology errors are categorized as follows: 1. Term Not Applied: A required glossary term was not used; a different or incorrect term was inserted instead, violating glossary rules. 2. Wrong Term: The glossary term appears but is used incorrectly in context—this includes typos, capitalization issues, plural/singular mismatches, or improper grammatical inflection.   Severity Levels Critical: Glossary misuse leads to confusion, misinterpretation, or disrupts key meaning. Major: The error affects tone or grammatical flow, though the meaning remains somewhat clear. Minor: Cosmetic or minor issues such as case sensitivity or small typos that don’t impact comprehension.   Example Source (English): S/N (Serial Number) Target (Russian): серийный номер (серийный номер)